A freight train takes 18 hours to travel the same distance that an express t rain travels in 15 hours. The rate of the express train is 15 miles per hour faster than the freight train. Find the rate at which each train travels.
Answer by macston(5194)

.
distance=rate x time
Since distance is the same:
rate of express x 15 hours=rate of freight x 18 hours
rate of express=rate of freight + 15mph
Let F=rate of freight train:
.
(F+15mph)(15hrs)=F(18hrs)
15F+225=18F
225=3F
75mph=F ANSWER 1: The freight train travels at 75 mph.
Express=Freight + 15 mph=75mph+15mph=90mph
ANSWER 2: The express train travels at 90 mph.
.
CHECK:
(90mph)(15hrs)=(75mph)(18hrs)
1350 miles=1350 miles